# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v1

@@ 7.1. Comparison with geometry
Recall that the isomorphism $\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})\cong N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ is constructed in [BravermanFinkelbergNakajima] by identifying the source with the homology of the affine Grassmannian and by using the derived Satake equivalence of [BezrukavnikovFinkelberg]. In this subsection, we compare the isomorphism of loc. cit. with the one constructed in this paper.
@@
Analogously to Section 5, one can define the negative version of the Kostant-Whittaker reduction
@@
for any Harish-Chandra bimodule $X$. Recall that the coefficients of the quantum minor $T^{1\ldots N}_{1\ldots N}(u)$ from (1.11) generate the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$. Let us identify it with the generating function of the equivariant Chern classes of the dual representation, see Remark 7.15:
@@
For every dominant weight $\lambda\in\Lambda$, the derived Satake equivalence of [BezrukavnikovFinkelberg] gives an isomorphism of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodules
@@
[…]
@@
Recall description of $\mathrm{Gr}^{\lambda}$ in terms of lattices, i.e. free $\mathbf{C}[[t]]$-modules of rank $N$ inside $\mathbf{C}((t))^{N}$ [Zhu]. We will only consider a simple case $\lambda=\omega_{1}^{*}$ (and $\lambda=\omega_{1}$ in Appendix B), where $\omega_{1}^{*}=(0,\ldots,0,-1)$. Then
@@
[…]
@@
$H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}^{*}},\underline{\mathbf{C}}_{\mathrm{Gr}^{\omega_{1}^{*}}})=H^{\bullet}_{\mathrm{GL}_{N}\times\mathbf{C}^{*}}(\check{\mathbb{P}}^{N-1})\cong\mathbf{C}[\hbar][A_{1},\ldots,A_{N}][c^{\mathrm{GL}_{N}}_{1}(\mathcal{S})]/(A(c^{\mathrm{GL}_{N}}_{1}(\mathcal{S}))).$
@@
By [BezrukavnikovFinkelberg, Lemma 4], the Kostant-Whittaker reduction functor is a full embedding when restricted to free Harish-Chandra bimodules. Therefore, we can lift $\phi$ to an automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$ which induces an automorphism of the parabolic restriction $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes\mathbf{C}^{N}/\!\!/N_{-}$ as of $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodules. According to the decomposition (4.2), it is isomorphic to a direct sum of $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodules
@@
where $\bar{v}_{i}$ are vectors (4.10). Since $\phi$ was defined on the non-generic bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$, the automorphism acts as $\bar{v}_{i}\mapsto f_{i}(w)\bar{v}_{i}$ for some polynomial function $f_{i}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})$ with a polynomial inverse, therefore, it should be constant $f_{i}(w)=c_{i}$. Moreover, since $\phi$ commutes with the diagonal action of the Weyl group $W=S_{N}\subset\mathrm{GL}_{N}$, we have $c_{i}=c_{j}$ for any $i,j$. Hence $\phi=c\cdot\mathrm{Id}$ by 4.5. ∎
@@
composed with Proposition 5.5 is given by $\Omega^{k}\mapsto(-1)^{N-1}c^{\mathrm{GL}_{N}}_{1}(\mathcal{S})^{k}$ for all $k$.
@@
The left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-module structures are compatible by (7.8). Let us show that this map intertwines the right actions. Following the proof of [CautisKamnitzer, Lemma 8.13], consider two short exact sequences on $\mathrm{Gr}^{\omega_{1}^{*}}$:
@@
which agrees with the formula of Proposition 4.7.
Therefore, this map is indeed an isomorphism of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodules such that $v_{1}\mapsto(-1)^{N-1}\in H^{\bullet}_{G}(\mathbb{P}^{N-1})$. The equivalence of Bezrukavnikov-Finkelberg gives an a priori different isomorphism with the same normalization condition by [FKPRW, Lemma 2.13]. However, it should coincide with the former by 7.12. ∎
Finally, recall the truncated shifted Yangian of 2.17. By [BravermanFinkelbergNakajima, Theorem B.18], there is a surjective homomorphism
@@
Here $[\mathrm{Gr}^{\omega_{1}}]\in H_{\bullet}^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr})$ is the fundamental class of the embedding $\mathrm{Gr}^{\omega_{1}}\hookrightarrow\mathrm{Gr}$, similarly for $[\mathrm{Gr}^{\omega_{1}^{*}}]$.
@@
[…]
is commutative. Here, the left vertical arrow is Theorem 7.9 and the lower horizontal one is the composition of (7.7) and the derived Satake equivalence (7.6).
@@
Observe that the composition of Theorem 5.9 and Proposition 7.13 gives an algebra map
@@
Likewise, the dual versions of Theorem B.3 and Proposition B.6 give
@@
By Theorem 7.9, the combination of algebraic maps gives an isomorphism
@@
The explicit identification (7.4) of [BezrukavnikovFinkelberg] comes from two facts. On the algebraic side, in loc. cit., the authors identify the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ with $W$-invariant functions on $\mathfrak{h}_{N}^{*}$ shifted by $-\hbar\rho_{N}$ via the projection
$\mathbf{Z}_{\mathfrak{gl}_{N}}\hookrightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\rightarrow\mathfrak{n}_{+}\backslash\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}),$
for instance, see [BezrukavnikovFinkelberg, Lemma 5]. By definition of the quantum minor (1.9), we have
@@ 7.2. Monopole operators
Recall the notations from Section 1.1. Let $\lambda$ be a dominant weight, and denote by $v^{*}_{\lambda}\in V_{\lambda}^{*}$ the vector dual to the lowest-weight vector in the representation $V_{\lambda}$. Observe that the function $\langle g\cdot v_{\lambda},v^{*}_{\lambda}\rangle$ is $N_{+}^{L}\times N_{+}^{R}$-invariant, hence defines an element in the Toda lattice $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$. The goal of this subsection is to compute these elements and to identify their images under the GKLO map for fundamental weights. This is an algebraic counterpart of the simplest case of the monopole operators formula [BravermanFinkelbergNakajima, (A.5)].
@@
Under the embedding ${}^{R}\iota_{N}$ of Example 3.5, we have
@@
[…]
@@
(recall that $\psi=1$). At the same time, it follows from Proposition 7.2 and (1.14) that
@@
[…]
where $e_{k}\colon(\mathbf{C}^{N})^{\otimes k}\rightarrow S^{k}\mathbf{C}^{N}$ is the symmetrizer map. Also, by Proposition 5.10, we have
@@
which, by Proposition 5.7, is given by the action of the element
@@
in the spherical subalgebra of $\mathfrak{H}^{\hbar}_{k}$ as in (2.7). By Proposition 2.7, it is equal to
@@
The next theorem computes the images of these elements under the GKLO map of Theorem 7.9.
@@
of Theorem 6.7. Recall that
@@
where $J$ runs over all multisubsets of $\{1,\ldots,N\}$ and $c_{J}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$ are some functions. In what follows, we will use Proposition 4.12 and notations thereof. By Proposition 7.16, we need to compute
@@
Therefore, the only non-zero indices in the sum are permutations of $\{1,\ldots,k\}$. Moreover, by $W$-invariance of Remark 7.10, it is enough to compute the coefficient $c_{\{1,\ldots,k\}}(w)$. By an easy induction on $k$, we have
@@
This is a rational analog of the simplest operator in the M-system in its shuffle realization, see [TsymbaliukGKLO, (104)] and [TsymbaliukGKLO, Proposition 6.13]. By [FinkelbergTsymbaliuk, Remark 9.19], its operators correspond to the $K$-theory classes of line bundles $\mathcal{O}(-k)$ restricted to the orbits $\mathrm{Gr}^{\omega_{i}}$ as in Section 7.1, where $\mathcal{O}(1)$ is the theta bundle on the affine Grassmannian. Therefore, Proposition 7.16 and Theorem 7.17 are in accordance with the derived Satake isomorphism of Theorem 7.14, since by [FKPRW, Lemma 2.6], we have $\langle g\cdot v_{\omega_{i}},v_{\omega_{i}}^{*}\rangle\mapsto[\mathrm{Gr}^{\omega_{i}}]$, up to a scalar factor. It would be interesting to find a rational analog of the full $M$-system.

### v2

@@ 7.1. Comparison with geometry
Recall that the isomorphism $\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})\cong N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ is constructed in [BFN19] by identifying the source with the homology of the affine Grassmannian and by using the derived Satake equivalence of [BF08]. In this subsection, we compare the isomorphism of loc. cit. with the one constructed in this paper.
@@
Analogously to Section˜5, one can define the negative version of the Kostant-Whittaker reduction
@@
for any Harish-Chandra bimodule $X$. Recall that the coefficients of the quantum minor $T^{1\ldots N}_{1\ldots N}(u)$ from (1.11) generate the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$. Let us identify it with the generating function of the equivariant Chern classes of the dual representation, see Remark˜7.15:
@@
For every dominant weight $\lambda\in\Lambda$, the derived Satake equivalence of [BF08] gives an isomorphism of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodules
@@
${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes V_{\lambda}/\!\!/_{\psi^{*}}\,N^{R}_{-}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\overline{\mathrm{Gr}^{\lambda}},\mathrm{IC}^{\lambda})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})}$
@@
Recall description of $\mathrm{Gr}^{\lambda}$ in terms of lattices, i.e. free $\mathbf{C}[[t]]$-modules of rank $N$ inside $\mathbf{C}((t))^{N}$ [Zhu17]. We will only consider a simple case $\lambda=\omega_{1}^{*}$ (and $\lambda=\omega_{1}$ in Appendix˜B), where $\omega_{1}^{*}=(0,\ldots,0,-1)$. Then
@@
(7.10) ${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varpi}_{N}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes(\mathbf{C}^{N})^{*}/\!\!/_{\psi^{*}}\,N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}^{*}},\underline{\mathbf{C}}_{\mathrm{Gr}^{\omega_{1}^{*}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})}$
@@
$H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}^{*}},\underline{\mathbf{C}}_{\mathrm{Gr}^{\omega_{1}^{*}}})=H^{\bullet}_{\mathrm{GL}_{N}\times\mathbf{C}^{*}}(\mathbb{P}^{N-1})\cong\mathbf{C}[\hbar][A_{1},\ldots,A_{N}][c^{\mathrm{GL}_{N}}_{1}(\mathcal{S})]/(A(c^{\mathrm{GL}_{N}}_{1}(\mathcal{S}))).$
@@
By [BF08, Lemma 4], the Kostant-Whittaker reduction functor is a full embedding when restricted to free Harish-Chandra bimodules. Therefore, we can lift $\phi$ to an automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$ which induces an automorphism of the parabolic restriction $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes\mathbf{C}^{N}/\!\!/N_{-}$ as of $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodules. According to the decomposition (4.2), it is isomorphic to a direct sum of $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodules
@@
where $\bar{v}_{i}$ are vectors (4.10). Since $\phi$ was defined on the non-generic bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$, the automorphism acts as $\bar{v}_{i}\mapsto f_{i}(w)\bar{v}_{i}$ for some polynomial function $f_{i}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})$ with a polynomial inverse, therefore, it should be constant $f_{i}(w)=c_{i}$. Moreover, since $\phi$ commutes with the diagonal action of the Weyl group $W=S_{N}\subset\mathrm{GL}_{N}$, we have $c_{i}=c_{j}$ for any $i,j$. Hence $\phi=c\cdot\mathrm{Id}$ by ˜4.5. ∎
@@
composed with Proposition˜5.5 is given by $\Omega^{k}\mapsto(-1)^{N-1}c^{\mathrm{GL}_{N}}_{1}(\mathcal{S})^{k}$ for all $k$.
@@
The left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-module structures are compatible by (7.8). Let us show that this map intertwines the right actions. Following the proof of [CK18, Lemma 8.13], consider two short exact sequences on $\mathrm{Gr}^{\omega_{1}^{*}}$:
@@
which agrees with the formula of Proposition˜4.7.
Therefore, this map is indeed an isomorphism of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodules such that $v_{1}\mapsto(-1)^{N-1}\in H^{\bullet}_{G}(\mathbb{P}^{N-1})$. The equivalence of Bezrukavnikov-Finkelberg gives an a priori different isomorphism with the same normalization condition by [Fin+18, Lemma 2.13]. However, it should coincide with the former by ˜7.12. ∎
Finally, recall the truncated shifted Yangian of ˜2.17. By [BFN19, Theorem B.18], there is a surjective homomorphism
@@
Here $\mathrm{Gr}^{\omega_{1}}]\in H_{\bullet}^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr})$ is the fundamental class of the embedding $\mathrm{Gr}^{\omega_{1}}\hookrightarrow\mathrm{Gr}$, similarly for $\mathrm{Gr}^{\omega_{1}^{*}}]$.
@@
${\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{\bullet}^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr})}$
is commutative. Here, the left vertical arrow is Theorem˜7.9 and the lower horizontal one is the composition of (7.7) and the derived Satake equivalence (7.6).
@@
Observe that the composition of Theorem˜5.9 and Proposition˜7.13 gives an algebra map
@@
Likewise, the dual versions of Theorem˜B.3 and Proposition˜B.6 give
@@
By Theorem˜7.9, the combination of algebraic maps gives an isomorphism
@@
The explicit identification (7.4) of [BF08] comes from two facts. On the algebraic side, in loc. cit., the authors identify the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$ with $W$-invariant functions on $\mathfrak{h}_{N}^{*}$ shifted by $\hbar\rho_{N}$ via the projection
[…]
for instance, see [BF08, Lemma 5]. By definition of the quantum minor (1.9), we have
@@ 7.2. Monopole operators
Recall the notations from Section˜1.1. Let $\lambda$ be a dominant weight, and denote by $v^{*}_{\lambda}\in V_{\lambda}^{*}$ the vector dual to the lowest-weight vector in the representation $V_{\lambda}$. Observe that the function $\langle g\cdot v_{\lambda},v^{*}_{\lambda}\rangle$ is $N_{+}^{L}\times N_{+}^{R}$-invariant, hence defines an element in the Toda lattice $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$. The goal of this subsection is to compute these elements and to identify their images under the GKLO map for fundamental weights. This is an algebraic counterpart of the simplest case of the monopole operators formula [BFN19, (A.5)].
@@
Under the embedding $^{R}\iota_{N}$ of Example˜3.5, we have
@@
$\rightarrow\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})}\ldots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{L})}\mathrm{D}_{\hbar}(\mathrm{GL}_{N})\xrightarrow{m}\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$
@@
(recall that $\psi=1$). At the same time, it follows from Proposition˜7.2 and (1.14) that
@@
${\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{}^{R}\iota_{N}\otimes\ldots\otimes{}^{R}\iota_{N}}$$\scriptstyle{e_{k}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})}$${\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\otimes S^{\bullet}\mathbf{C}^{N}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{}^{R}\iota_{N}}$
where $e_{k}\colon(\mathbf{C}^{N})^{\otimes k}\rightarrow S^{k}\mathbf{C}^{N}$ is the symmetrizer map. Also, by Proposition˜5.10, we have
@@
which, by Proposition˜5.7, is given by the action of the element
@@
in the spherical subalgebra of $\mathfrak{H}^{\hbar}_{k}$ as in (2.7). By Proposition˜2.7, it is equal to
@@
The next theorem computes the images of these elements under the GKLO map of Theorem˜7.9.
@@
of Theorem˜6.7. Recall that
@@
where $J$ runs over all multisubsets of $\{1,\ldots,N\}$ and $c_{J}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$ are some functions. In what follows, we will use Proposition˜4.12 and notations thereof. By Proposition˜7.16, we need to compute
@@
Therefore, the only non-zero indices in the sum are permutations of $\{1,\ldots,k\}$. Moreover, by $W$-invariance of Remark˜7.10, it is enough to compute the coefficient $c_{\{1,\ldots,k\}}(w)$. By an easy induction on $k$, we have
@@
This is a rational analog of the simplest operator in the M-system in its shuffle realization, see [Tsy23, (104)] and [Tsy23, Proposition 6.13]. By [FT19, Remark 9.19], its operators correspond to the $K$-theory classes of line bundles $\mathcal{O}(-k)$ restricted to the orbits $\mathrm{Gr}^{\omega_{i}}$ as in Section˜7.1, where $\mathcal{O}(1)$ is the theta bundle on the affine Grassmannian. Therefore, Proposition˜7.16 and Theorem˜7.17 are in accordance with the derived Satake isomorphism of Theorem˜7.14, since by [Fin+18, Lemma 2.6], we have $\langle g\cdot v_{\omega_{i}},v_{\omega_{i}}^{*}\rangle\mapsto[\mathrm{Gr}^{\omega_{i}}]$, up to a scalar factor. It would be interesting to find a rational analog of the full $M$-system.